OpenStudy (anonymous):

Can anyone sole the IVP y'' + 2ky' + ( (k^2) + (w^2))y = 0, y(0) = 1, y'(0) = -k

OpenStudy (anonymous):

Wow, that's a hairy one. Are k and w constants?

OpenStudy (anonymous):

usually with this sort of IVP you'd use laplace iirc

OpenStudy (anonymous):

yes they are constants

OpenStudy (anonymous):

you hev to use homogeneous linear ODEs to solve it

OpenStudy (dumbcow):

step 1: set up characteristic equation \[r^{2} +2kr + (k^{2}+w^{2})=0\] \[r = -k \pm wi\] complex roots, so use Eulers Formula to put in sin/cos form general solution \[y(t) = c_{1}e^{-kt}\cos(wt) + c_{2}e^{-kt}\sin(wt)\] apply initial conditions y(0) = 1 --> c1 = 1 \[y'(t) = kwe^{-kt}\sin(wt) - kwc_{2}e^{-kt}\cos(wt)\] y'(0) = -k --> -kwc2 = -k --> c2 = 1/w \[y(t) = e^{-kt} \cos(wt) + \frac{1}{w}e^{-kt} \sin(wt)\]